Weekly summary, July 09-15

Rick Kessler (kessler@hep.uchicago.edu)
Fri, 16 Jul 1999 16:55:20 -0500

Summary of Week: July 09-15 Runs: 13839-13915

Hours of beam used/delivered : 105.1 / 118.2 = 0.89

Vac00 (week) : 425 K (64.2 K/day)
Vac00 (99run) : 1.37 M (46.6 K/day)

Live DAQ with beam/elapsed time (week) : 0.69
Live DAQ with beam/elapsed time (99run) : 0.69

Comment: main loss due to 28 hour Tevatron shutdown;
Small KTEV losses are spread out over many systems,
no "killer" incident this week.

The HA shaper/splitter was bypassed so we have no monitoring
of individual channels; we can only monitor HA scalers.

hours of beam:
Shift used/delivered comment
---------------------------------------------------------------------
Fri July 09 OWL: 1.3/2.0
DAY: 5.3/5.5
EVE: 6.3/6.4 Switch from 4E12 ==> 7E12

Sat July 10 OWL: 5.0/6.0 Installed new run control ==> bad idea!
DAY 5.5/6.5 Replace fuse in HA Shaper/splitter
EVE 6.5/8.0 DPMT swap!

Sun July 11 OWL: 5.6/6.1 DPMT swap continued from previous shift
DAY: 7.7/8.5 Minor DA problems
EVE: 7.0/8.0 L3 monitor takes x2 CPU ==> DAQDEAD time

Mon July 12 OWL: 5.7/7.0 Pipeline problems (1 hour)
DAY: 6.7/8.2 DLT drive failure; false fire alarm
EVE: 7.4/8.4 Mu-run destination error from bad DLT drive

Tue July 13 OWL: 6.4/6.5
DAY: ------- SHUTDOWN: new RC to label tape in parallel
EVE: ------- SHUTDOWN: HA SS Hell begins ...

Wed July 14 OWL: ------- SHUTDOWN: HA splitter bypassed.
DAY: 2.5/3.4 Access to re-adjust BA & HA DC offsets.
EVE: 6.4/6.5 One L3 monitor CPU died.

Thu July 15 OWL: 6.0/6.3 Pipeline crate re-booted. SEM > 7.5E12
DAY: 7.0.7.3
EVE: 6.8/7.6
---------------------------------------------------------------------
TOTAL 105.1/118.2
---------------------------------------------------------------------